Block 𝐴 has a weight 𝑀𝐴 and block 𝐡 has a weight 𝑀𝐡. Once block 𝐡 is set into downward motion, it descends at a constant speed. a) Calculate the coefficient of kinetic friction between block 𝐴 and the tabletop. b) A cat, also of weight 𝑀𝐴, falls asleep on top of block 𝐴. If block 𝐡 is now set into downward motion, what is its acceleration (magnitude and direction)?Β 


1
Expert's answer
2021-10-10T15:54:33-0400

(a) If the blocks move at constant speed, the net force is zero, and forces acting on Block B are


"0=W_B-T=0."

Forces on block A are:


"T-\mu W_A=0,\\\space\\
\mu=\frac{T}{W_A}=\frac{W_B}{W_A}."

(b) When there is a cat on top of block A, the force of friction increases (because the normal force increases proportionally to the combined weight of block A and weight of the cat), while the weight of block B (and, thus, the tension in the cord) remains constant, thus, the acceleration is zero.